# Hierarchical DLO Routing with Reinforcement Learning and In-Context Vision-language Models

### v1

@@ IV-D Real-world Robot Experiments
We conducted real robot long-horizon routing tasks using a Franka Emika Panda robot. A wrist-mounted, calibrated Intel RealSense D415 camera captures top-down images of the tabletop scene at a resolution of $1280\times 720$. The MoveIt planner [34] was used to generate continuous trajectories based on the motion primitives predicted by our model or predefined motion skills. The overall system was built using the Robot Operating System (ROS) framework [35], and clips and DLO are segmented using SAM2 [36] with an NVIDIA 5090 GPU. The reference points for DLO segmentation are initially set around the fixture and iteratively added during segmentation. Importantly, our policies did not require additional data for fine-tuning or domain adaptation during real robot testing. We compare our approach to the best-performing baseline in the 4 tasks mentioned in Tab. II and report the success rate and average number of clips inserted as shown in Tab. III. The routing process is shown in Fig. 5. Although the success rate in real experiments is lower than in simulation, our method still achieves strong performance, reaching 62.5% success despite challenges from calibration errors, perception noise, and sim-to-real transfer. In contrast, the Fixed Order baseline performs significantly worse, highlighting that real-world routing requires explicit failure reasoning and closed-loop decision making to handle unexpected variations and disturbances. The common failure cases in simulation and in the real world are summarized in Fig. 6. Faulty planning has been observed in a real experiment where the robot executes the ’flatten’ command even though the DLO is already inserted into the clip, leading to a grasping failure and a collision with the clip. Another suboptimal plan observed in real and simulation is that the planner ignores the emerging DLO segment in the zoomed-in image, incorrectly concludes that insertion is incomplete, and issues another insertion action. While this behavior does not necessarily result in a collision since our RL policy can select an appropriate grasping location, it still unnecessarily prolongs the routing process and leads to a suboptimal plan.
@@
Success Rate (%) $\uparrow$ 37.5 62.5
Avg. #Clips Ins. $\uparrow$ 1.75 2.625
@@ V Limitations & Future Work
While our long-horizon cable routing framework provides a promising solution for DLO routing in diverse scenes, several failure cases remain unsolved. Addressing faulty or suboptimal planning is an important direction for future work. One potential approach is to fine-tune the VLM planner to improve planning stability. Another direction is to integrate new skills into the VLM planner automatically. This process could be automated by providing the planner with sequences of action images and letting it reason through the usage and definition of skills. Such an approach would enable large-scale skill integration and improve generalization to increasingly complex deformable manipulation tasks.
@@ VI Conclusion
In this work, we presented a hierarchical framework for long-horizon cable routing that integrates a high-level VLM planner with a key low-level RL skill. The VLM leverages in-context reasoning and failure recovery to generate reliable task plans and skill selections, while the RL insertion policy robustly handles complex DLO dynamics. Through extensive simulation and real-world experiments, we demonstrated that our approach outperforms fixed-order and heuristic baselines, achieving strong generalization to 4-clip settings. These results highlight the importance of combining high-level reasoning with task-aware low-level control for scalable DLO manipulation.

### v2

@@ IV-D Real-world Robot Experiments
We conducted real robot long-horizon routing tasks using a Franka Emika Panda robot. A wrist-mounted, calibrated Intel RealSense D415 camera captures top-down images of the tabletop scene at a resolution of $1280\times 720$. The MoveIt planner [3] was used to generate continuous trajectories based on the motion primitives predicted by our model or predefined motion skills. The overall system was built using the Robot Operating System (ROS) framework [26], and clips and DLO are segmented using SAM2 [27] with an NVIDIA 5090 GPU. The reference points for DLO segmentation are initially set around the fixture and iteratively added during segmentation. Our RL policy uses state-based observations extracted from segmentation. During simulation training, we randomize rope configurations to capture a wide distribution of states, enabling the policy to generalize to real-world rope configurations. Consequently, our policy did not require additional data for fine-tuning or domain adaptation during real robot testing. We evaluate our method on 8 clip configurations and compare our approach to the Fixed Order baseline across the 4 tasks mentioned in Tab. II. We report the success rate and the average number of clips inserted, as shown in Tab. III. The routing process is shown in Fig. 5. Although the success rate in real experiments is lower than in simulation, our method still achieves strong performance, reaching 62.5% success despite challenges from calibration errors, perception noise, and sim-to-real transfer. In contrast, the Fixed Order baseline performs significantly worse, highlighting that real-world routing requires explicit failure reasoning and closed-loop decision making to handle unexpected variations and disturbances. The common failure cases in simulation and in the real world are summarized in Fig. 6. Faulty planning has been observed in a real experiment where the robot executes the ’flatten’ command even though the DLO is already inserted into the clip, leading to a grasping failure and a collision with the clip. Another suboptimal plan observed in real and simulation is that the planner ignores the emerging DLO segment in the zoomed-in image, incorrectly concludes that insertion is incomplete, and issues another insertion action. While this behavior does not necessarily result in a collision since our RL policy can select an appropriate grasping location, it still unnecessarily prolongs the routing process and leads to a suboptimal plan.
@@
SR $\uparrow$ 37.5 62.5
Clips $\uparrow$ 1.75 2.625
@@ V Limitations & Future Work
While our long-horizon DLO routing framework provides a promising solution for DLO routing in diverse scenes, several failure cases remain unsolved. Addressing faulty or suboptimal planning is an important direction for future work. One potential approach is to fine-tune the VLM planner to improve planning stability. Another direction is to integrate new skills into the VLM planner automatically. This process could be automated by providing the planner with sequences of action images and letting it reason through the usage and definition of skills. Such an approach would enable large-scale skill integration and improve generalization to increasingly complex deformable manipulation tasks.
@@ VI Conclusion
In this work, we presented a hierarchical framework for long-horizon DLO routing that integrates a high-level VLM planner with a key low-level RL skill. The VLM leverages in-context reasoning and failure recovery to generate reliable task plans and skill selections, while the RL insertion policy robustly handles complex DLO dynamics. Through extensive simulation and real-world experiments, we demonstrated that our approach outperforms other baselines, achieving strong generalization to 4/5-clip settings. This highlights the importance of combining high-level reasoning with task-aware low-level control for scalable DLO manipulation.